OpenStudy (rissa):

Write an equation in point-slope form for the line through the given point with the given slope. (4, 3); m = 5/3

OpenStudy (asnaseer):

The equation of a line with slope m and passing through points (x1, y1) can be written as:\[y=y_1+ m(x-x_1)\]

OpenStudy (amistre64):

almost, barring for typos thats correct lol

OpenStudy (amistre64):

\[y-y_1= m(x-x_1)\]

OpenStudy (asnaseer):

where is my typo?

OpenStudy (amistre64):

unless "=" means "-" and "+" means "=" ... it fine

OpenStudy (asnaseer):

what you wrote is just a rearrangement of what I wrote - I'm lost...

OpenStudy (amistre64):

i know that; but the arrangement is whats key here point slope is defined by the arrangement

OpenStudy (asnaseer):

I am not really /for/ remembering things by route - I think it is more important to understand the principals.

OpenStudy (amistre64):

me too, but convention is convention :)

OpenStudy (rissa):

so y - 3 = 5/3 (x-4) ?

OpenStudy (asnaseer):

and remember - I said "can be written as..."

OpenStudy (amistre64):

yes, thats good

OpenStudy (asnaseer):

@risa - correct

OpenStudy (amistre64):

the equation of a line can be "written" in many different ways; but only a few f them have specific names applied to them ...

OpenStudy (amistre64):

my favorites tend to be intercept form and an amalgamation of point slope and slope intercept: y = mx-mPx+Py